OpenStudy (anonymous):

Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = (x-4)2-8 Shift the graph of y = x2 right 4 units and then up 8 units. Shift the graph of y = x2 right 4 units and then down 8 units. Shift the graph of y = x2 left 4 units and then down 8 units. Shift the graph of y = x2 down 4 units and then left 8 units.

OpenStudy (anonymous):

Shift the graph of y = x2 right 4 units and then down 8 units.

OpenStudy (anonymous):

and i also needed help with Describe how the graph of y = x2 can be transformed to the graph of the given equation. y = (x+17)2 Shift the graph of y = x2 left 17 units. Shift the graph of y = x2 down 17 units. Shift the graph of y = x2 up 17 units. Shift the graph of y = x2 right 17 units.

OpenStudy (anonymous):

Shift the graph of y = x2 left 17 units.

OpenStudy (anonymous):

Please understand HOW to do this though. Would you like an explanation?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Basically this equation gives you 2 things: y= (x+a) + b "a" will give you the horizontal transformation. (right and left) If it is a negative number, it will be translated to the right. If it is positive, it will be translated to the left. "b" gives you the vertical transformation.(up and down). If it is positive, it is up If it is negative, it is down

OpenStudy (anonymous):

THANK YOU Its better now